OpenStudy (anonymous):

What would I do in this situation with exponents?: I have 5x^4(2x^3-3x^2+5x-1) I know that in this case I would multiply the exponents like 4(2) but then once I do that, I end up with 10^12-15x^8+5x^2-5x. What's next?

jimthompson5910 (jim_thompson5910):

5x^4*2x^3 = (5*2)*(x^4*x^3) = 10x^(4+3) = 10x^7 So it's actually 10x^7 and NOT 10x^12 You do the same for the other terms

jimthompson5910 (jim_thompson5910):

remember to multiply 5x^4 with every term in the parenthesis

OpenStudy (anonymous):

so, for a problem like (x^2+4x+2)(x+7) would equal to 11x^2+30x+14

jimthompson5910 (jim_thompson5910):

no, right off the bat I know it should be a cubic since x^2 times x = x^3

OpenStudy (anonymous):

Can you please explain that problem with steps? I'm having serious trouble with exponents :/

jimthompson5910 (jim_thompson5910):

(x^2+4x+2)(x+7) (x+7)(x^2+4x+2) ... rearrange the terms x(x^2+4x+2) + 7(x^2+4x+2) ... expand x*x^2 + x*4x + x*2 + 7x^2 + 7*4x + 7*2 ... distribute x^3 + 4x^2 + 2x + 7x^2 + 28x + 14 .. multiply x^3 + 11x^2 + 30x + 14 ... Combine like terms. So (x^2+4x+2)(x+7) expands and simplifies to x^3 + 11x^2 + 30x + 14

OpenStudy (anonymous):

I got x^3+13x^2+42. I'm doing something wrong

jimthompson5910 (jim_thompson5910):

look through those steps and you should see where you went wrong

OpenStudy (anonymous):

I noticed I'm mostly having trouble with the rearranging and the whole rules of what to do with exponents when adding and subtracting. I know multiplying and dividing I add them,

OpenStudy (anonymous):

I know now. Thank you!!!!!!!!

jimthompson5910 (jim_thompson5910):

yw